# On uniqueness of KP soliton structures

### v1

@@ 6.1. Proof in the mZK case
In this section we provide proofs for Theorem 1.17 for the cubic versión ZK model (1.22). Notice that, as above, this equation can be recast in terms of mKdV (B.9) as follows:
$0=\left(-4u_{t}+\partial_{x_{1}}^{3}u+6u^{2}\partial_{x_{1}}u\right)+\partial_{x_{1}}(\Delta_{c}u),$
where $\Delta_{c}u=\sum_{j=2}^{d}\partial_{x_{j}}^{2}u$. For simplicity in the notation, let $x=x_{1}$. Replacing $u=2\partial_{x}F(\Theta)$ in (6.1), and integrating once in $x$, we obtain
@@
$+\sum_{j=2}^{d}\left(\Theta_{x_{j}x_{j}}F^{\prime}(\Theta)+\Theta_{x_{j}}^{2}F^{\prime\prime}(\Theta)\right)_{x}$
Notice in contrast of ZK case, the setting $Ai(\Theta)=0$, $W(\Theta)=0$ with $W_{x_{i}}^{F}=0$ produces a contradiction in (6.2), when considering $F=2\arctan$. Thus the condition $\Lambda_{x_{j}}^{F}(\Theta)=\Theta_{x_{j}x_{j}}F^{\prime}(\Theta)+\Theta_{x_{j}}^{2}F^{\prime\prime}(\Theta)=0$ enters in action. Similarly as in the ZK case, we arrive to the equation
@@
$+\Theta_{x}^{2}\left(h^{\prime\prime}+\frac{3}{s}h^{\prime}+8h^{3}\right)\left(\Theta\right)+\sum_{j=2}^{d}\left(\Lambda_{x_{j}}^{F}(\Theta)\right)_{x},$
where $h(s)=\frac{1}{2}F^{\prime}(s)$. Now we finish the proof. If $u$ is a mKdV soliton (see (B.11)), it is easy to check that (6.3) is satisfied. Now, if
$Ai(\Theta)=W(\Theta)=\Lambda_{x_{j}}^{F}(\Theta)=0,\quad j=2,\ldots,d,$ (6.4)
@@
$\Lambda_{x_{j}}^{F}(\Theta)=\Lambda_{x_{j}}^{2\arctan}(\Theta)=0,\quad i=2,\ldots,d.$
Now, from $W(\Theta)=0$ we get
$\Theta=a(x_{2},\ldots,x_{d},t)\exp\left(xb(x_{2},\ldots,x_{d},t)\right)+c(x_{2},\ldots,x_{d},t).$
Also since $\Lambda_{x_{j}}^{2\arctan}(\Theta)=0$ we will get that $a,b,c$ are only time dependent. Indeed if $a:=a(x_{2},\ldots,x_{d},t)$, $b:=b(x_{2},\ldots,x_{d},t)$, $c:=c(x_{2},\ldots,x_{d},t)$, replacing $\Theta$ in $\Lambda_{x_{j}}^{F}(\Theta)=0$ for each $j=2,\ldots,d$, one has
$0=$ $A_{0}+A_{1}\exp(bx)+A_{2}x\exp(bx)+A_{3}x^{2}\exp(bx)$
$+B_{1}\exp(2bx)+B_{2}x\exp(2bx)+B_{3}x^{2}\exp(2bx)$
$+C_{1}\exp(3bx)+C_{2}x\exp(3bx)+C_{3}x^{2}\exp(3bx),$
@@
Since $C_{3}=a^{3}b_{x_{j}}^{2}$ need to be zero, we have $b$ is constant in $x_{j}$. From $C_{2}=0$, taking in mind $w=a^{-1}$ as a change of variable of this ODE, we conclude that
$a(x_{j})=\frac{1}{Ax_{j}^{2}+Bx_{j}+C}.$ (6.5)
@@ 2. $\left(ii\right)$
$W\left(\Theta\right)=\Theta_{xx}^{2}-\Theta_{x}\Theta_{xxx}=\hbox{Wr}\left(\Theta_{x},\Theta_{xx}\right),$
@@ 1. $\left(i\right)$
[…]
@@ B.2. The mKdV case
The first term contains nothing but $W_{x}\left(\Theta\right):=\Theta_{xx}-\frac{\Theta_{x}^{2}}{\Theta}$, which will be required to be zero. Notice that $\Theta_{k}$ in (B.11) does satisfy this condition, i.e., $W_{x}(\Theta_{k})=0$. In the opposite direction, $W_{x}\left(\Theta\right)=0$ implies $\Theta\left(\frac{\Theta_{x}}{\Theta}\right)_{x}=0$, leading to $\Theta\left(t,x\right)=\exp\left(k\left(t\right)x+b\left(t\right)\right)$. From the Airy condition $\Theta_{xxx}-4{\Theta}_{t}=0$ one gets
@@
Finally, the second term in (B.13) (being zero now) can be written as the radial solution $h=h\left(s\right)$, $s=\left|x\right|>0$ of $\Delta h+8h^{3}=0$ in dimension 4. Requiring positive solutions (by hypothesis, $h\left(s\right):=\frac{1}{2}F^{\prime}\left(s\right)>0$), these are in $H^{1}\left(\mathbb{R}^{4}\right)$ and by classical Talenti-Aubin arguments, one has $h\left(s\right)=\frac{1}{1+s^{2}}$, giving, finally, $F=2\arctan$ if we assume $F\left(0\right)=0$. We conclude the following result:
@@
valid now in the full region. One can easily see that for any $\alpha,\beta>0$, $\delta=\frac{1}{4}(3\beta^{2}-\alpha^{2})$, and $\gamma=\frac{1}{4}(\beta^{2}-3\alpha^{2})$, one has that
$\Theta=\pm\frac{\beta}{\alpha}\frac{\sin(\alpha(x+\delta t))}{\cosh(\beta(x+\gamma t))},$
@@
$\Theta=\frac{e^{\sqrt{c_{1}}(x+c_{1}t)}+e^{\sqrt{c_{2}}(x+c_{2}t)}}{1-\rho^{2}e^{\sqrt{c_{1}}(x+c_{1}t)+\sqrt{c_{2}}(x+c_{2}t)}},\quad\rho:=\frac{\sqrt{c_{1}}-\sqrt{c_{2}}}{\sqrt{c_{1}}+\sqrt{c_{2}}},$

### v2

@@ 6.1. Proof in the mZK case
In this section we provide proofs for Theorem 1.17 for the cubic version ZK model (1.22). Notice that, as above, this equation can be recast in terms of mKdV (B.9) as follows:
$0=\left(-4u_{t}+\partial_{x_{1}}^{3}u+6u^{2}\partial_{x_{1}}u\right)+\partial_{x_{1}}\left(\Delta_{c}u\right),$
where $\Delta_{c}u=\sum_{j=2}^{d}\partial_{x_{j}}^{2}u$. For simplicity in the notation, let $x=x_{1}$. Replacing $u=2\partial_{x}F\left(\Theta\right)$ in (6.1), and integrating once in $x$, we obtain
@@
$+\sum_{j=2}^{d}\left(\Theta_{x_{j}x_{j}}F^{\prime}\left(\Theta\right)+\Theta_{x_{j}}^{2}F^{\prime\prime}\left(\Theta\right)\right)_{x}$
Notice in contrast of ZK case, the setting $Ai\left(\Theta\right)=0$, $W\left(\Theta\right)=0$ with $W_{x_{i}}^{F}=0$ produces a contradiction in (6.2), when considering $F=2\arctan$. Thus the condition $\Lambda_{x_{j}}^{F}\left(\Theta\right)=\Theta_{x_{j}x_{j}}F^{\prime}\left(\Theta\right)+\Theta_{x_{j}}^{2}F^{\prime\prime}\left(\Theta\right)=0$ enters in action. Similarly as in the ZK case, we arrive to the equation
@@
$+\Theta_{x}^{2}\left(h^{\prime\prime}+\frac{3}{s}h^{\prime}+8h^{3}\right)\left(\Theta\right)+\sum_{j=2}^{d}\left(\Lambda_{x_{j}}^{F}\left(\Theta\right)\right)_{x},$
where $h\left(s\right)=\frac{1}{2}F^{\prime}\left(s\right)$. Now we finish the proof. If $u$ is a mKdV soliton (see (B.11)), it is easy to check that (6.3) is satisfied. Now, if
$Ai\left(\Theta\right)=W\left(\Theta\right)=\Lambda_{x_{j}}^{F}\left(\Theta\right)=0,\quad j=2,\ldots,d,$ (6.4)
@@
$\Lambda_{x_{j}}^{F}\left(\Theta\right)=\Lambda_{x_{j}}^{2\arctan}\left(\Theta\right)=0,\quad i=2,\ldots,d.$
Now, from $W\left(\Theta\right)=0$ we get
$\Theta=a\left(x_{2},\ldots,x_{d},t\right)\exp\left(xb\left(x_{2},\ldots,x_{d},t\right)\right)+c\left(x_{2},\ldots,x_{d},t\right).$
Also since $\Lambda_{x_{j}}^{2\arctan}\left(\Theta\right)=0$ we will get that $a,b,c$ are only time dependent. Indeed if $a:=a\left(x_{2},\ldots,x_{d},t\right)$, $b:=b\left(x_{2},\ldots,x_{d},t\right)$, $c:=c\left(x_{2},\ldots,x_{d},t\right)$, replacing $\Theta$ in $\Lambda_{x_{j}}^{F}\left(\Theta\right)=0$ for each $j=2,\ldots,d$, one has
$0=$ $A_{0}+A_{1}\exp\left(bx\right)+A_{2}x\exp\left(bx\right)+A_{3}x^{2}\exp\left(bx\right)$
$+B_{1}\exp\left(2bx\right)+B_{2}x\exp\left(2bx\right)+B_{3}x^{2}\exp\left(2bx\right)$
$+C_{1}\exp\left(3bx\right)+C_{2}x\exp\left(3bx\right)+C_{3}x^{2}\exp\left(3bx\right),$
@@
Since $C_{3}=a^{3}b_{x_{j}}^{2}$ needs to be zero, we have $b$ is constant in $x_{j}$. From $C_{2}=0$, taking in mind $w=a^{-1}$ as a change of variable of this ODE, we conclude that
$a\left(x_{j}\right)=\frac{1}{Ax_{j}^{2}+Bx_{j}+C}.$ (6.5)
@@ 2. $\left(ii\right)$
$W\left(\Theta\right)=\Theta_{xx}^{2}-\Theta_{x}\Theta_{xxx}=\hbox{Wr}\left(\Theta_{xx},\Theta_{x}\right),$
@@ 1. $\left(i\right)$
    $Ai\left(\Theta\right)=c_{0}\exp\left(F\left(\Theta\right)\right)-3W\left(\Theta\right)+3F\left(\Theta\right)\int_{0}^{x}\exp\left(-F\left(\Theta\right)\right)\left(W\left(\Theta\right)\right)_{x}\left(t,s\right)\mathrm{d}s.$ (B.8)
@@ B.2. The mKdV case
The first term contains nothing but $W_{x}\left(\Theta\right):=\Theta_{xx}-\frac{\Theta_{x}^{2}}{\Theta}$, which will be required to be zero. Notice that $\Theta_{k}$ in (B.11) does satisfy this condition, i.e., $W_{x}\left(\Theta_{k}\right)=0$. In the opposite direction, $W_{x}\left(\Theta\right)=0$ implies $\Theta\left(\frac{\Theta_{x}}{\Theta}\right)_{x}=0$, leading to $\Theta\left(t,x\right)=\exp\left(k\left(t\right)x+b\left(t\right)\right)$. From the Airy condition $\Theta_{xxx}-4{\Theta}_{t}=0$ one gets
@@
Finally, the second term in (B.13) (being zero now) can be written as the radial solution $h=h\left(s\right)$, $s=\left|x\right|>0$ of the elliptic, nonlinear PDE $\Delta h+8h^{3}=0$ in dimension 4. Requiring positive solutions (by hypothesis, $h\left(s\right):=\frac{1}{2}F^{\prime}\left(s\right)>0$), these are in $H^{1}\left(\mathbb{R}^{4}\right)$ and by classical Talenti-Aubin arguments, one has $h\left(s\right)=\frac{1}{1+s^{2}}$, giving, finally, $F=2\arctan$ if we assume $F\left(0\right)=0$. We conclude the following result:
@@
valid now in the full region. One can easily see that for any $\alpha,\beta>0$, $\delta=\frac{1}{4}\left(3\beta^{2}-\alpha^{2}\right)$, and $\gamma=\frac{1}{4}\left(\beta^{2}-3\alpha^{2}\right)$, one has that
$\Theta=\pm\frac{\beta}{\alpha}\frac{\sin\left(\alpha\left(x+\delta t\right)\right)}{\cosh\left(\beta\left(x+\gamma t\right)\right)},$
@@
$\Theta=\frac{e^{\sqrt{c_{1}}\left(x+c_{1}t\right)}+e^{\sqrt{c_{2}}\left(x+c_{2}t\right)}}{1-\rho^{2}e^{\sqrt{c_{1}}\left(x+c_{1}t\right)+\sqrt{c_{2}}\left(x+c_{2}t\right)}},\quad\rho:=\frac{\sqrt{c_{1}}-\sqrt{c_{2}}}{\sqrt{c_{1}}+\sqrt{c_{2}}},$